OpenStudy (anonymous):

Determine an equivalent expression for sin 3x cos x + cos 3x sin x . A. 4 sin x B. 2 sin x cos x C. 4 sin x cos x D. 2 sin 2x cos 2x

OpenStudy (anonymous):

D. Just substitute x=1 in to the expressions and they should give the same as the expression.

myininaya (myininaya):

sin3x cox+cos3x sinx =sin(3x+x) =sin(4x) =sin(2x+2x) =sin2x cos 2x+sin2x cos 2x =2 sin2x cos 2x